A helium filled balloon has a volume of 2.75 dm3 at 20.0°C. The volume of the balloon changes to 2.07 dm3 when placed outside on a cold day. What is the temperature outside in °C?


1
Expert's answer
2022-04-07T04:05:55-0400

V1/T1 = V2/T2

2.75 L / (20.0 + 273.15) K = 2.07 L / T2

T2 = 220.66 K = (220.66 - 273.15)°C = -52.5°C